# Self-generated electrokinetic flows from active-charged boundary patterns

### v1

@@ Abstract
We develop a hydrodynamic description of self-generated electrolyte flow in capillaries whose bounding walls feature both non-uniform distributions of charge and non-uniform active ionic fluxes. The hydrodynamic velocity arising in such a system has components that are forbidden by symmetry in the absence of charge and fluxes. We show that flow states with unidirectional and circulatory components emerge when modulated boundary patterns are offset by a selection of flux-charge phase differences associated with the spatial mismatch between sites of peak flux and charge densities on the wall. Self-induced local ionic elevation and depletion, constantly disrupting a non-uniform double layer, promotes directed gradients yielding persistent body forces that generate longitudinal fluid motion along the capillary. Our work quantifies a boundary-driven mechanism based on active-charged patterns that produce self-sustained electrolyte flow in confined environments and exists even in the absence of any external bulk-imposed fields or gradients. It provides a theoretical framework for understanding the effect of ionic boundary properties that are relevant in biological or soft matter systems and can be utilized in nanofluidics and iontronics.
@@ I Introduction
Self-sustained fluid motion is an essential mechanism involved in various cellular processes that regulate intracellular transport [1, 2, 3, 4, 5]. Micro- and nano-fluidic technologies inspired by biological or soft matter systems have witnessed rapid development in recent years. This includes the emerging field of iontronics which employs ions as charge carriers, for instance, in neuromorphic devices that can be comprised of fluidic channels or capillaries analogous to the axon of neurons [6, 7, 8, 9, 10]. The directed flow of charged fluids such as an ionic solution or electrolyte in confined environments has thus become an important area of investigation. In particular, the underlying physical parameters and confinement boundary properties that control and mediate such flows need to be understood and clarified [11, 12, 13, 14, 15].
@@
A well-known mechanism for micro- and nano-scale transport of electrolyte near charged interfaces employs gradients of electric potential or ionic concentration. These are types of electrokinetic flows referred to as electro-osmosis or ionic diffusio-osmosis, respectively. On one hand, this can be achieved through externally applied electric fields or by imposing bulk ionic gradients. Mobile ions and fixed surface charges are brought into balance by forming an electric double layer (EDL) at the interface, and the applied field or gradient exerts body forces acting within this layer where the ions are dislocated along with the surrounding viscous liquid, causing a net fluid motion in response [16, 17, 18, 19, 20]. On the other hand, electrokinetic flows can alternatively be induced through surface-generated ionic fluxes that are spatially nonuniform or asymmetric [21, 22, 23, 24]. Variations in fluxes along the surface lead to local ionic gradients, which in turn induce fluid flow. Ionic fluxes naturally occur at active sites such as membrane ion pumps in a biological context [25, 26, 27, 28], or can be chemically produced via surface reactants grafted in catalytic micro-channels or capillaries [29, 30, 31]. Furthermore, in the realm of self-propelling active colloids, it has been shown that charged Janus particles with asymmetric fluxes of ions on their surface move via ionic self-phoresis [32, 33, 34, 35, 36].
@@
Spatial variations of interfacial properties such as surface charge are also known to affect electrokinetic flow fields and net velocity [37, 38, 39]. In the case of externally driven elecroosmosis, both theoretical [40, 41, 42, 43] and experimental [44, 45, 46] works have shown that surface charge variation leads to different flow patterns. Heterogeneous surface charge triggers additional gradients on top of the constant applied field, and induce nonuniform flows in the bulk such as recirculating and multidirectional flows when alternating charges are patterned periodically on channel walls. This feature can be utilized in fluidic devices that are funtionalized for mixing of ionic solutions. On the other hand, when charges travel on the wall of a capillary in the form of a wave [47, 48, 49, 50], a non-vanishing unidirectional (zero mode) velocity is generated, carrying downstream the liquid electrolyte [51]. In light of these past and current developments, the role of boundary heterogeneities in electrolyte transport is becoming more clear and has opened new avenues for technological applications.
@@
Figure 2: Flow states produced by flux and charge boundary modulations. (a) The $z$-component velocity profile shapes for circulatory ±$\pm W$, unidirectional $U$, and intermediate $U\pm W$. The stream plots of steady state fluid velocity in (r,z) obtained analytically from Eq. (29) and (30) discussed in Sec. IV for case of zero anionic flux $j_{0-}=0$ and same relative flux-charge peak amplitudes $\delta_{+}=\delta_{\sigma}$ in Eq. (18), at different flux-charge phase difference $\beta$ (with $\alpha_{+}=0$): leading to (b) fully circulatory flow at $\beta=\pi$ and (c) unidirectional dominant flow at $\beta=\pi/2$, and (d) intermediate flow at $\beta=3\pi/4$. The corresponding modulation forms of the cationic (+) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction, and the color map indicates the velocity magnitude given in terms of the characteristic flow speed $v_{0}$. We show here a capillary with radius $R=10\lambda$ and modulation wavelength $l=200\lambda$ given in terms of the Debye length $\lambda$.
@@
In this work, we investigate the flow of electrolyte confined by walls endowed with spatially-varying charge and wall ionic activity. We consider a cylindrical capillary with periodically patterned surface charge and fluxes of ions, as illustrated in Fig. 1. We show that the hydrodynamic velocity contains terms which exist only when the aforementioned boundary mechanisms are simultaneously present. The significance of these terms is that they describe a physical effect which is absent in systems lacking these boundary features. In particular, we establish the existence of a zero mode velocity, that is, a non-vanishing unidirectional velocity component which is self-generated along the capillary. This effect arises from a selective spatial mismatch at the boundary between sites of peak flux and charge densities that breaks translation-reflection symmetry. The effect is due to constant disruption of a tightly bound double layer leading to strong lateral body forces driving the bulk viscous flow. Estimates of the effect at various electrolyte concentrations and conditions show that it is of considerable magnitude and thus expected be observed with great facility in an experiment. We emphasize that this is a surface-driven effect and exists even in the absence of any external bulk-imposed fields and concentration gradients.
@@
We employ the continuum framework of the Poisson-Nernst-Planck-Navier-Stokes (PNPNS) equations, and carry out theoretical analytic approximations and finite-element numerical simulations. We show that boundary-modulated charge and active ionic fluxes self-generate two basic flow state components, a circulatory and a unidirectional (zero mode) along the capillary length, as shown in panels (b) and (c) of Fig. 2, respectively. These flow states critically depend on the phase differences between ionic fluxes and charge modulation patterns and are determined by a number of selection rules. Because the double layer is non-uniform in the longitudinal direction of the capillary, due to the varying wall charge coupled with strong enough ionic fluxes, there is significant ionic elevation and depletion, constantly knocking the EDL off its equilibrium state. Consequently, this leads to self-induced ionic gradients which are in-phase or out-of phase with respect to the commensurate electric field. In the latter case the electric body force is such as to drive the circulation of a sequence of adjacent vortices. In the former case the electric body force is always of a single sign and drives the electrolyte in a single direction. Moreover, we find a variation of ion diffusivity or relative flux amplitude can lead to velocity reversal and switching of the flow state. Our work illustrates a boundary-driven mechanism that controls the transport of confined electrolytes based on non-uniform patterning of active and charged sites.
@@
This article has been organized as follows. In Sec. II we establish the form the velocity components acquire in the presence of charge and flux amplitudes as these are allowed by symmetry. Sec. III describes the continuum PNPNS theory which we employ to obtain the flow patterns in the small-field and long wavelength approximations. In Sec. IV we employ numerical simulations to illustrate and discuss the different flow states generated for special cases of ionic activity and study their dependence on phase difference of the flux and charge modulation. These conclusions are corroborated by the analytical approximations developed in section  III. We quantify and provide estimates of the net unidirectional flow in Sec. V showing that it can reach considerable magnitudes. Likewise, in sec. VI we employ analytical approximations and numerical simulations to demonstrate the effect of ionic diffusivity mismatch. Finally, we conclude by summarizing our main results and giving an outlook on future directions in Sec. VII. The Appendices include details of our analytical calculations and details of the finite-element numerical techniques.
@@ II Symmetry of self-generated flow
The hydrodynamic velocity of the electrolyte arising in such a confinement has radial and axial components described by $\mathbf{v}=v_{r}(r,z)\hat{{r}}+v_{z}(r,z)\hat{{z}}$ in the axisymmetric cylindrical coordinates $(r,z)$ as these are displayed in figure 1. Since the velocity is a polar vector and changes sign under spatial inversion, in the presence of both boundary effects, the following contributions to the axial component are allowed
@@
where $m,m^{\prime}\in\{+,-\}$ and summation is implied over repeated indices. The coefficients $\gamma$ above are odd functions of $k$. Thus, under spatial inversion the velocity changes sign. We show below that the first term in the right hand side of (3) gives rise to circulatory behavior and the second term leads a zero mode which unidirectionally drives the electrolyte along the capillary, for example shown in panels (b) and (c) of Fig. 2, respectively.
@@ III Continuum hydrodynamic theory
The characteristic length scale associated with the size of the EDL is the Debye length given by $\lambda=\sqrt{\varepsilon k_{B}T/e^{2}C_{\infty}}$, where $\varepsilon$ is the electric permittivity and $k_{B}T$ is the characteristic thermal energy of liquid water at room temperature. Here $C_{\infty}$ is the total bulk ion concentration such that both ionic species approach the same value $C_{\infty}/2$ at the reservoirs. We consider ionic concentrations in the range of $C_{\infty}\sim 10^{-2}-10^{2}$ mM and the confinement sizes in the regime where this Debye length is non-vaninshing, in that it is not negligible with respect to capillary radius like $R\sim\lambda-10^{2}\lambda$. Here we study the behavior at long time scales such that fluxes of the ions and surface charge relaxes to steady (time-independent) values. More specifically, we are looking at the system after some space-charge relaxation time $\tau>R\lambda/D_{\pm}$ [52, 53], where $D_{\pm}$ is the diffusivity of the ion species which typically $\sim 10^{-9}-10^{-8}$ m2 sec-1. In the following subsections we detail the governing equations of our steady state model, the small-field approximations employed to solve the system, and the analytical solutions obtained in the long wavelength mode.
@@ III.1 Governing equations and boundary conditions
The ion concentrations considered are dilute such that the solution is modeled as an ideal mixture, and the Péclet number Pe $\sim v_{0}\lambda/D_{\pm}$ is small in our regime of length scales for the characteristic flow speed $v_{0}$. Thus, the net (volume) flux of the ionic species are
@@
where the more dominant contributions are provided by the diffusive flux and the electrochemical flux, while the advective contribution, subject to conditions of the aforementioned discussion, is not taken into account. Thus, the conservation of mass $\nabla\cdot\mathbf{J}_{\pm}=0$ gives the steady state Nernst-Planck equations,
@@
where $\hat{n}$ is the unit vector poining into the liquid and $\sigma$ was defined in (2). Second, is the spatially varying cation/anion ($\pm$) fluxes coming inwards (from the wall into the fluid domain) and going outwards (from the fluid domain into the wall),
@@
where the $j_{\pm}$ were defined in (1). Additionally, on a solid wall we impose the no-slip boundary condition $\mathbf{v}=\mathbf{0}$. For the purposes of our analytical formulation given below, the capillary is approximated to be of infinte extent in the $z$-direction.
III.2 Small-field perturbations
@@ III.3 Self-induced body force
$\begin{split}\tilde{F}_{z}=\frac{\tilde{k}}{2}I\big{(}\sqrt{1+\tilde{k}^{2}}\tilde{r}\big{)}\ I(\tilde{k}\tilde{r})\bigg{[}h_{0}&+h_{c}\cos(2\tilde{k}\tilde{z})\\ &+h_{s}\sin(2\tilde{k}\tilde{z}))\bigg{]}\,,\end{split}$ (20)
@@
The body force expression (20) clarifies the character of the resultant flow. The first term in the square brackets of Eq. (20) leads to a force always pointing in the same direction and gives rise to a zero mode (unidirectional) velocity, as it will be discussed below. The second and third terms in the square brackets of Eq. (20) lead to the formation of vortices at every half-period of the wall flux/charge variation. The symmetry properties of the resulting velocity field discussed in Eq. (3) are also evident in Eq. (20). Under spatial inversion the force changes sign. Thus, it gives rise to velocity fields that will also change sign as predicted in Eq. (3).
III.4 Long wavelength approximation
To obtain some limiting analytical insight into the effect we continue by performing the long wavelength approximation to the govering fields. The velocity field $\mathbf{v}=(v_{r},0,v_{z})$ can be expressed with respect to the (dimensionless) Stokes streamfunction $\psi(r,z)$ in the form

### v2

@@ Abstract
We develop a hydrodynamic description of self-generated electrolyte flow in capillaries whose bounding walls feature both non-uniform distributions of charge and non-uniform active ionic fluxes. The hydrodynamic velocity arising in such a system has components that are forbidden by symmetry in the absence of charge and fluxes. However, when these two boundary mechanisms are simultaneously present, they can lead to a symmetry broken state where steady flows with both unidirectional and circulatory components emerge. We show that these flow states arise when modulated boundary patterns of charge and fluxes are offset by a flux-charge phase difference, which is associated with the separation between sites of their peak densities on the wall. Mismatch in diffusivity of cationic and anionic species can modify the flow states and becomes an enhancing factor when fluxes of both ion species are being produced together at the same site. We demonstrate that this mechanism can be realized with a microfluidic generator which is powered by enzyme-coated patches that catalyzes reactants in the solution to produce fluxes of ions. The local ionic elevation or depletion that disrupts a non-uniform double layer, promotes self-induced gradients yielding persistent body forces to generate bulk fluid motion. Our work quantifies a boundary-driven mechanism behind self-sustained electrolyte flow in confined environments that exists without any external bulk-imposed fields or gradients. It provides a theoretical framework for understanding the combined effect of active and charged boundaries that are relevant in biological or soft matter systems, and can be utilized in electrofluidic and iontronic applications.
@@ I Introduction
Self-sustained fluid motion is an essential mechanism involved in various cellular processes that regulate intracellular transport [1, 2, 3, 4, 5]. Micro- and nano-fluidic technologies inspired by biological or soft matter systems have witnessed rapid development in recent years. This includes the emerging field of iontronics which employs ions as charge carriers, for instance, in neuromorphic devices that can be comprised of fluidic chambers analogous to the axon of neurons [6, 7, 8, 9, 10]. The directed flow of charged fluids such as an ionic solution or electrolyte in confined environments has thus become an important area of investigation. In particular, the underlying physical parameters and confinement boundary properties that control and mediate such flows need to be understood and clarified [11, 12, 13, 14, 15, 16].
@@
A well-known mechanism for micro- and nano-scale transport of electrolyte near charged interfaces employs gradients of electric potential or ionic concentration. These are types of electrokinetic flows referred to as electro-osmosis or ionic diffusio-osmosis, respectively. On one hand, this can be achieved through externally applied electric fields or by imposing bulk ionic gradients. Mobile ions and fixed surface charges are brought into balance by forming an electric double layer (EDL) at the interface, and the applied field or gradient exerts body forces acting within this layer where the ions are dislocated along with the surrounding viscous liquid, causing a net fluid motion in response [17, 18, 19, 20, 21]. On the other hand, electrokinetic flows can alternatively be induced through surface-generated ionic fluxes that are spatially nonuniform or asymmetric [22, 23, 24, 25]. Variations in fluxes along the surface lead to local ionic gradients, which in turn induce fluid flow. Ionic fluxes can naturally occur at active sites such as membrane ion pumps in a biological context [26, 27, 28, 29], or can be chemically produced via surface reactants grafted in catalytic micro-channels or capillaries [30, 31, 32, 33, 34]. Furthermore, in the realm of self-propelling active colloids, it has been shown that charged Janus particles with asymmetric fluxes of ions on their surface move via ionic self-phoresis [35, 36, 37, 38, 39].
@@
Spatial variations of interfacial properties such as surface charge are also known to affect electrokinetic flow fields and net velocity [40, 41, 42]. In the case of externally driven elecroosmosis, both theoretical [43, 44, 45, 46] and experimental [47, 48, 49, 50] works have shown that surface charge variation leads to different flow patterns. Heterogeneous surface charge triggers additional gradients on top of the constant applied field, and induce nonuniform flows in the bulk such as recirculating and multidirectional flows when alternating charges are patterned periodically on channel walls. This feature can be utilized in fluidic devices that are functionalized for mixing and pumping of ionic solutions.
@@
Figure 2: Flow states produced by flux and charge boundary modulations. (a) The $z$-component velocity profile shapes for circulatory ±$\pm W$, unidirectional $U$, and intermediate $U\pm W$. The stream plots of steady state fluid velocity in $(\tilde{r},\tilde{z})$ obtained analytically from Eq. (29) and (30) discussed in Sec. IV for case of zero anionic flux $j_{0-}=0$ and same relative flux-charge peak amplitudes $\delta_{+}=\delta_{\sigma}$ in Eq. (18), at different flux-charge phase difference $\beta$ (with $\alpha_{+}=0$): leading to (b) fully circulatory flow at $\beta=\pi$ and (c) unidirectional dominant flow at $\beta=\pi/2$, and (d) intermediate flow at $\beta=3\pi/4$. The corresponding modulation forms of the cationic (+) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction, and the color map indicates the velocity magnitude. We show here in dimensionless units of $\tilde{r}=r/\lambda$, $\tilde{z}=z/\lambda$ and $\bar{v}=v/\lambda$ s-1 for a capillary with radius $R=10\lambda$ and modulation wavelength $l=200\lambda$ given in terms of the Debye length $\lambda$.
@@
Whereas unidirectional electrolyte flow induced by time-dependent boundary voltages has been realized both experimentally [51, 52] and theoretically [53, 54, 55, 56], a time-independent concept that would replace these approaches and circumvent early challenges, is still elusive. In this paper we show that the hydrodynamic velocity of an electrolyte confined by walls endowed with time-independent spatially-varying charge and ionic activity (Fig. 1), contains terms which exist only when the aforementioned boundary mechanisms are simultaneously present. The significance of these terms is that they describe a physical effect which is absent in systems lacking these boundary features. We emphasize that this is a boundary-driven effect that exists without any external bulk-imposed electric fields and ionic concentration gradients. In particular, we establish the existence of a zero mode velocity, that is, a non-vanishing unidirectional velocity component that is self-generated in conjunction with a circulatory component along the capillary. Disruption of a non-uniform double layer via local ionic elevation or depletion, leads to strong body forces driving the bulk viscous flow. We find that this effect arises from a flux-charge phase difference associated with spatial separation between sites of peak flux and charge densities on the boundary. Moreover, a mismatch in diffusivity between cationic and anionic species modifies the flow states and becomes an enhancing factor when fluxes of both cations and anions are being produced together at the same site. Estimates of the effect at various electrolyte concentrations and conditions show that it is of considerable magnitude and thus expected be observed with great facility in an experiment. We further demonstrate how this flux-charge mechanism can be realized in a concrete example of a microfluidic generator powered by chemical reactions with enzyme-coated wall patches.
@@
We employ the continuum framework of the Poisson-Nernst-Planck-Navier-Stokes (PNPNS) equations, and carry out theoretical analytic approximations and finite-element numerical simulations. We show that boundary-modulated charge and active ionic fluxes self-generate two basic flow state components, a circulatory and a unidirectional (zero mode) along the capillary length, as shown in panels (b) and (c) of Fig. 2, respectively. These flow states depend on the phase differences between ionic fluxes and charge modulation patterns and are determined by a number of selection rules. Because the EDL is non-uniform in the longitudinal direction of the capillary, due to the varying wall charge coupled with strong ionic fluxes, there is significant ionic elevation and depletion, constantly knocking it off a uniform equilibrium state. Consequently, this leads to self-induced ionic gradients which are in-phase or out-of phase with respect to the commensurate electric field. In the latter case the electric body force is such as to drive the circulation of a sequence of adjacent vortices. In the former case the electric body force is always of a single sign and drives the electrolyte in a single direction. We also find that the switching of cationic-anionic diffusivity can lead to velocity reversal and change of the flow state.
The mechanism of flux-charge separation and diffusivity mismatch can be realized with an enzyme powered microfluidic generator [30, 33, 37, 34]. In particular, we demonstrate this using a microfluidic chamber as in Fig. 1 where an urease-coated patch [33, 37] is grafted in the capillary wall which catalyzes reactants in the liquid solution releasing a localized in-flux of ions as end products. Together with this flux patch, we also consider a separate wall patch constructed with silica [37] or polymer coating [47, 49] which carries some local surface charge. We show that such a pair of patches generates unidirectional and circulatory flows with localized vortex patterns and the flow speed depends on the flux-charge patch separation distance. Furthermore, we compare two cases of previous works on urease patches [33, 37] where the pair of cation-anion end products are different with distinct diffusivity mismatch, and find that the flow speed is enhanced by the larger mismatch. Taken together, our work describes a boundary-driven mechanism based on the combined effect of non-uniform patterning of active and charged sites that controls the transport of confined electrolytes.
@@
This article has been organized as follows. In Sec. II we establish the form the velocity components acquire in the presence of charge and flux amplitudes as these are allowed by symmetry. Sec. III describes the continuum PNPNS theory which we employ to obtain the flow patterns in the small-amplitude and long-wavelength approximations. In Sec. IV we employ numerical simulations to illustrate and discuss the different flow states generated for special cases of ionic activity and study their dependence on phase difference of the flux and charge modulation. These conclusions are corroborated by the analytical approximations developed in section  III. We quantify and provide estimates of the net unidirectional flow in Sec. V showing that it can reach considerable magnitudes. Likewise, in sec. VI we employ analytical approximations and numerical simulations to demonstrate the effect of ionic diffusivity mismatch. Then we present in Sec. VII a concrete example of an enzyme powered microfluidic generator. Finally, we conclude by summarizing our main results and giving an outlook on future directions in Sec. VIII. The Appendices include details of our analytical calculations and details of the finite-element numerical techniques.
@@ II Symmetry of self-generated flow
The hydrodynamic velocity of the electrolyte arising in such a confinement has radial and axial components described by $\mathbf{v}=v_{r}(r,z)\hat{{r}}+v_{z}(r,z)\hat{{z}}$ in the axisymmetric cylindrical coordinates $(r,z)$ as these are displayed in figure 1. In the presence of both boundary effects, the following contributions to the axial component are allowed
@@
where $m,m^{\prime}\in\{+,-\}$ and summation is implied over repeated indices. The coefficients $\gamma$ above are odd functions of $k$. Since the velocity is a polar vector, it changes sign under spatial and domain inversion $(r,z)\rightarrow(-r,-z)$ and $k\rightarrow-k$. We show below that the first term in the right hand side of (3) gives rise to circulatory behavior and the second term leads a zero mode which unidirectionally drives the electrolyte along the capillary, for example shown in panels (b) and (c) of Fig. 2, respectively.
@@ III Continuum hydrodynamic theory
The characteristic length scale associated with the size of the EDL is the Debye length given by $\lambda=\sqrt{\varepsilon k_{B}T/e^{2}C_{\infty}}$, where $\varepsilon$ is the electric permittivity and $k_{B}T$ is the characteristic thermal energy of liquid water at room temperature. Here $C_{\infty}$ is the total bulk ion concentration such that both ionic species approach the same value $C_{\infty}/2$ at the reservoirs. We consider bulk ionic concentrations in the range of $C_{\infty}\sim 10^{-2}-10^{2}$ mM corresponding to Debye lengths $\lambda\sim 1-10^{2}$ nm, which is non-vanishing in that it is small but not negligible with respect to the confinement size like capillary radii $R\sim 0.1-10$ $\mu$m. Here we study the behavior at long time scales such that fluxes of the ions and surface charge relaxes to steady (time-independent) values. More specifically, we are looking at the system after some space-charge relaxation time $\tau>R\lambda/D_{\pm}$ [57, 58], where $D_{\pm}$ is the diffusivity of the ion species which typically $\sim 10^{-9}-10^{-8}$ m2 s-1. In the following subsections we detail the governing equations of our steady state model, the analytical approximations employed to solve the system, and solutions obtained.
@@ III.1 Governing equations and boundary conditions
The ion concentrations considered are dilute such that the solution is modeled as an ideal mixture, and the Péclet number Pe $\sim v_{0}\lambda/D_{\pm}$ is small in our regime of length scales for the characteristic flow speed $v_{0}$. Thus, the net (volume) fluxes of the ionic species are
@@
where the dominant contributions are provided by the diffusive flux and the electrochemical flux, while the advective contribution, subject to conditions of the aforementioned discussion, is not taken into account. Thus, the conservation of mass $\nabla\cdot\mathbf{J}_{\pm}=0$ gives the steady state Nernst-Planck equations,
@@
where $\hat{n}$ is the unit vector pointing into the liquid and $\sigma$ was defined in (2). Second, is the spatially varying cation/anion ($\pm$) fluxes coming inwards (from the wall into the fluid domain) and going outwards (from the fluid domain into the wall),
@@
where the $j_{\pm}$ were defined in (1). Additionally, on a solid wall we impose the no-slip and no-penetration boundary condition $\mathbf{v}=\mathbf{0}$. For the purposes of our analytical formulation given below, the capillary is approximated to be of infinte extent in the $z$-direction.
III.2 Small field perturbations
@@ III.3 Self-induced body force
$\begin{split}\tilde{f}_{z}\approx\frac{\tilde{k}}{2}I\big{(}\sqrt{1+\tilde{k}^{2}}\tilde{r}\big{)}\ I(\tilde{k}\tilde{r})\bigg{[}h_{0}&+h_{c}\cos(2\tilde{k}\tilde{z})\\ &+h_{s}\sin(2\tilde{k}\tilde{z}))\bigg{]}\,,\end{split}$ (20)
@@
The body force expression (20) clarifies the character of the resultant flow. The first term in the square brackets of Eq. (20) leads to a force always pointing in the same direction and gives rise to a zero mode (unidirectional) velocity, as it will be discussed below. The second and third terms in the square brackets of Eq. (20) lead to the formation of vortices at every half-period of the wall flux and charge variations. The symmetry properties of the resulting velocity field discussed in Sec. II are also evident in Eq. (20), that is, under the spatial and domain inversion this body force changes sign. Thus, it gives rise to velocity fields that will also change sign as predicted in Eq. (3).
III.4 Long wave mode
To obtain some analytical insight into the effect we continue by performing the long-wavelength approximation to the governing fields. The velocity field $\mathbf{v}=(v_{r},0,v_{z})$ can be expressed with respect to the (dimensionless) Stokes streamfunction $\psi(r,z)$ in the form